OpenStudy (anonymous):

average rate of change for f(x) x=4 to x=7?

OpenStudy (triciaal):

i think you forgot f(x)

OpenStudy (anonymous):

its a whole assignment so i dont know what to do like i converted f(x) to a vertex form then into standard but i dont know

OpenStudy (triciaal):

rate of change is the slope (Y2 -y1)/(x2 - x1 ) this is also the derivative depends on what class you are in dy/dx

OpenStudy (anonymous):

f(x) = 2 (x - 6)^2 + 3 is f(x) in vertex i think

OpenStudy (anonymous):

then i got y = 2x^2 - 24x - 69 for standard form

OpenStudy (kirbykirby):

avg rate of chage = \[\frac{f(7)-f(4)}{7-4} \]

OpenStudy (anonymous):

but then wouldnt that just be the same thing on top and bottom? f(3) / 3?

OpenStudy (kirbykirby):

\(f(7)-f(4)\ne f(7-4)\)

OpenStudy (kirbykirby):

and \(f(7)\ne 7\)

OpenStudy (anonymous):

im so confused

OpenStudy (kirbykirby):

If f(x) = 2 (x - 6)^2 + 3 How do you evaluate f(7) ?

OpenStudy (anonymous):

2( 7 - 6)^2 + 3 ?

OpenStudy (kirbykirby):

yes

OpenStudy (kirbykirby):

do the same with f(4),

OpenStudy (anonymous):

and then divide ?

OpenStudy (kirbykirby):

well subtract those two results, and then divide by (7-4) The idea here is that the rate of change is basically taking two points on your parabola, and then drawing a line through those points. Then the average rate of change is basically the slope of that lie you drew. Now, the slope of a line can be expressed as \[\frac{y_2-y_1}{x_2-x_1} \] , and remember you can write \(y=f(x)\), so that the formula becomes \[ \frac{f(x_2)-f(x_1)}{x_2-x_1}\]

OpenStudy (anonymous):

for f(4) i got 19 and for f(7) i got 7

OpenStudy (anonymous):

so i divide 19 and 7 by 3?

OpenStudy (kirbykirby):

hm not quite :S I'm not sure how you calculated that. Remember the rules for exponents and parenthesis: Parenthesis first, then exponents, then multiplication/division, then addition/subtraction

OpenStudy (anonymous):

2 (4-6)^2 + 3 2 (-2)^2 +3 -4 ^2 +3 16 + 3 = 19

OpenStudy (anonymous):

2 (7-6) ^2 + 3 2 (1)^2 +3 2^2 + 3 4 + 3 = 7

OpenStudy (kirbykirby):

2 (-2)^2 +3 this becomes 2(4)+3 since (-2)^2= 4 and it's an exponent, so you do it before the multiplication

OpenStudy (anonymous):

11?

OpenStudy (anonymous):

so 11 and 19?

OpenStudy (kirbykirby):

2 (4-6)^2 + 3 = 11 -------------------------- 2(7 - 6)^2 + 3 2(1)^2 + 3 2(1)+3 =5

OpenStudy (anonymous):

oh i see both of my mistkes.. then what do we do with 11 and 5?

OpenStudy (kirbykirby):

so from the formula \[\frac{f(7)-f(4)}{7-4}=\frac{5-11}{7-4}=\frac{-6}{3}=-2 \]

OpenStudy (anonymous):

thank you so much! i was so lost before i wanted to break windows in my house

OpenStudy (anonymous):

so average rate of change is -2?

OpenStudy (kirbykirby):

yup

OpenStudy (kirbykirby):

yw :), lol don't break any windows!